OpenStudy (anonymous):

Calculate the speed of an 8.4 × 10 4 kg airliner with a kinetic energy of 1.4 × 10 9 J. Answer in units of m/s

OpenStudy (a_clan):

1 J= 1 (kg-m^2)/s^2 1.4 × 10 J = 1.4 × 10^9 (kg-m^2)/s^2 KE = (1/2).m.v^2 Substitute m=8.4 × 10^4 kg and KE = 1.4 × 10^9 (kg-m^2)/s^2